A lady went to abank with rupees 1000 . She asked the cashier to give her rupees 10 and 50 currency notes in return she got 40 currency notes in all . Find the number of each kind of currency notes please help me to slove this problem i have eaxam today
Answer by ikleyn(52787)

.
Let T be the number of 10-rs notes.

Then the number of 50-rs notes is (40-T).


The money equation is

10*T + (40-T)*50 = 1000.


Simplify and solve for T:

10T + 2000 - 50T = 1000  ====>  -40T = 1000-2000 = -1000  ====>  T = {{(-1000}/(-40)}}} = 25.


Answer.  25 10-rs notes and 40-25 = 15 50-rs notes.


Check.   25*10 + 15*50 = 250 + 750 = 1000.    ! Correct !.